# Tenspiler: A Verified Lifting-Based Compiler for Tensor Operations

### v1

@@ Abstract
To enable developers in leveraging such DSLs while preserving their current programming paradigm, we describe our experience in building Tenspiler, a verified lifting-based compiler that uses program synthesis to translate sequential programs written in general-purpose programming languages (e.g., C++ or Python code that does not leverage any specialized framework or accelerator) into tensor operations. Central to Tenspiler is our carefully crafted yet simple intermediate language, named Tensir, that expresses tensor operations. Tensir enables efficient lifting, verification, and code generation. Unlike classical pattern-matching-based compilers, Tenspiler uses program synthesis to translate input code into Tensir, which is then compiled to the target API / ISA. Currently, Tenspiler already supports six
DSLs, spanning a broad spectrum of software and hardware environments. Furthermore, we show that new backends can be easily supported by Tenspiler by adding simple pattern-matching rules for Tensir. Using 10 real-world code benchmark suites, our experimental evaluation shows that by translating code to be executed on 6 different software frameworks and hardware devices, Tenspiler offers on average 105$\times$ kernel and 9.65$\times$ end-to-end execution time improvement over the fully-optimized sequential implementation of the same benchmarks.
@@ 1 Introduction
In this paper, we describe our experience in building a tensor compiler that addresses these challenges. We introduce Tenspiler—a compiler designed to automate the transpilation of code to multiple tensor processing frameworks and hardware accelerators. Tenspiler uses verified lifting [12] (VL), a technique using inductive program synthesis to infer provably equivalent program summaries expressed using a user-defined intermediate representation (IR), and generate executable code from the synthesized summary to the target DSL. In contrast to conventional compilers that rely on pattern-matching to compile code, VL uses a search-based technique for the translation process. The two key steps of VL are:
@@
  • Verification: Once lifted, a theorem prover is used to validate if the synthesized summary is functionally equivalent to the input. If so, executable code is produced by calling the user-provided code generator from the summary, otherwise, another summary is generated by the search phase.
The key to making lifting efficient lies in the design of the IR (i.e., how the target DSL is modeled). In prior work [5, 6, 7], each function or instruction exposed by the target DSL is modeled explicitly. While doing so makes the search efficient, such explicit modeling makes the compiler hard to extend to other DSLs. With Tenspiler, we introduce, for the first time, a single unified IR, Tensir, that is designed for tensor operations and can easily generate code to multiple tensor processing software frameworks and hardware accelerators. Surprisingly, Tensir is a small language based on tensor algebra that includes commonly used vector and matrix operations. While other unifying IR exists (e.g., MLIR [24]), they are targeted for classical pattern-matching compilers. As we will discuss in Sec. 5 and Section 4.2, Tensir instead is designed for synthesis-based compilers and thus aims to enable both efficient search and verification.
In summary, this experience paper makes the following contributions:
@@
    We implemented Tenspiler, a verified lifting-driven transpiler built using Tensir as the modeling language. We demonstrate the effectiveness of Tenspiler by using it to lift real-world code from 10 different suites to 6 different open-source and commercially-available tensor processing software frameworks and hardware accelerators. We demonstrate the ease of constructing such transpilers by building one for MLX, a new tensor processing framework that was released only four months ago, using less than 200 lines of code in Sec. 6.
@@ 2 Overview
Figure 1: End-to-End example of using Tenspiler to transpile code
Tenspiler takes in C/C++ or Python code as input [^1] Tenspiler currently supports a subset of the C/C++ and Python language (in particular it does not support code that uses pointers or objects, which we have not encountered such use in our benchmarks). It also expects any external libraries used in the input to be functionally modeled, which is how Tenspiler currently supports code that uses the STL::vector library. and transpiles it to a functionally equivalent program that leverages different software frameworks and hardware accelerators (details described in  Section 4.3) for tensor computation. As mentioned, Tenspiler uses verified lifting to first translate the source program into Tensir. Unlike traditional pattern-matching compilers, Tenspiler formulates code translation as a search for a program expressed in Tensir that is provably semantic-equivalent to the input. Doing so avoids the need to devise pattern-matching rules and prove their correctness. To make the search scalable, instead of directly searching within the DSL exposed by each target, we designed a high-level IR called Tensir that abstracts away the low-level implementation details of each DSL operator and captures only their semantics, unifying various DSLs into a common set of tensor operators.
Tenspiler uses a program synthesizer (currently Rosette [34], a synthesizer for finite domain theories) to lift the input code to Tensir during the search phase. The synthesized output is then verified using a theorem prover (currently, an SMT solver, CVC5 [10]) for the unbounded domain. “Unbounded domain” means the verification is performed for all possible program states, not just a bounded set of states (e.g., all states where integers are represented using 8 bits) that Rosette considers during the synthesis phase.
Once verified, Tenspiler then translates the Tensir program to the concrete syntax of the target DSL using simple pattern-matching rules.
@@
and is defined recursively on each element. Meanwhile, t_s
performs element-wise scalar operations on tensor x using the scalar value a and is similarly defined. Importantly, both operators are purely functional models of the tensor operations that lack implementation details that a specific target might leverage (e.g., tiling, vectorization, etc). The idea is that if S can be expressed using only these operators via lifting, then the lifted program can be easily translated to the targeted backends.
@@
Tenspiler uses symbolic search to solve the synthesis problem. Symbolic search is typically implemented through enumerative or deductive search, and using constraint-solving approaches which often rely on domain-specific heuristics to scale.
As a SyGuS problem, symbolic search is implemented as enumerating different expressions over a user-provided grammar, where the grammar encodes all possible combinations of operators in the target DSL up to a specified depth. However, as the depth increases, the number of choices grows exponentially, making the search intractable
As we will discuss in Sec. 5, Tensir is designed to make synthesis scalable. For S, the synthesis phase returns the following solution:
@@ 3 The Tensir Intermediate Representation
Comparison with MLIR. MLIR is a compiler infrastructure that enables the representation and transformation of code at various levels of abstraction. The core idea behind MLIR is to provide a unified IR that can capture the semantics of the program at different levels of detail (dialects), from high-level abstractions down to low-level, target-specific instructions. Developers can use MLIR by progressively lowering the code through different dialects until it reaches a level suitable for the target hardware. While MLIR and its dialects offer a powerful infrastructure for progressively targeting multiple hardware backends, we found that the existing dialects do not fully support all the operators required for our use case. Independently, both the linalg and tensor dialects do not support all the operators Tensir supports. For example, the $select$ operator, which is crucial for image processing kernels that apply operations conditioned on pixel values, is not supported by any MLIR dialects. Additionally, unifying these dialects can be challenging for developers. Instead of unifying, recent work such as mlirSynth [14] has explored using program synthesis to translate between different MLIR dialects. In contrast, Tensir is designed to be flexible and easily extensible. Developers can add new operators to Tensir by simply describing their high-level semantics, and new backend support can be incorporated by defining simple pattern-matching rules. This approach allows developers to extend Tensir without going into the intricacies of MLIR. Moreover, Tensir can practically be compiled into different MLIR dialects, providing developers the flexibility to leverage the MLIR infrastructure if desired.
@@ 3.1 Language Definition
$a\in accessOp$ ::=:= $take(t,l)\;\mid\;tail(t,l)\;\mid\;slice(t,d,l_{1},l_{2})$
$l$ ::=:= $Integer\;Literal\;\mid\;size(t)\;\mid\;s_{comp}$
@@
Figure 2: Tensir grammar
The recursive nature of Tensir’s grammar allows tensor operations to be composed, facilitating the expression of complex algorithms encountered in source code. Tensir extends its expressiveness beyond tensor operations by also including a control flow operator (controlflowOp)
It integrates control flow through the ite operator, enabling conditional logic into tensor computation. This operator is crucial for translating real-world loopy programs that contain branches.
@@
  • slice(t, d, s, e) extracts a contiguous sub-tensor from t from indices s (inclusive) to e (exclusive) along dimension d.
  • size(t) returns the total number of elements in tensor t.
@@ 4 Transpiling Code Using Tenspiler
As shown in Fig. 3, Tenspiler is designed to translate a program in high-level languages, source (S),
into another program that leverages different tensor processing infrastructures. Tenspiler currently support a vector processing library (NumPy) for CPU execution, DL frameworks (PyTorch, TensorFlow, MLX) for GPU execution, and ISAs for specialized hardware accelerators (Gaudi, Gemmini). Tenspiler is a verified-lifting-based compiler, leveraging search to find a program within the target domain. Instead of relying on traditional pattern-matching rules, Tenspiler translates source programs with a 3-phase workflow: 1. synthesis, 2. verification, and 3. backend code generation. Tenspiler uses a single IR, Tensir, to facilitate all 3 phases of its workflow. Tensir is designed to include tensor processing operators common to all target backends. As shown in the figure, the synthesis phase takes in S and generates a program summary expressed using Tensir. Then, in the verification phase, Tenspiler verifies the generated summaries to ensure their semantic equivalence with S. Finally, in the code generation phase, the Tensir program is translated to the concrete syntax of the target DSL(s).
@@ 4.1 Synthesis
For Tenspiler, the specification is to find a functionally equivalent program to S. Various methods exist to express this specification, such as using input-output examples, bounded-model checking, and verification conditions (VC) [19]. In Tenspiler, we use VCs as the specification for the synthesis phase as it provides full guarantees (i.e., for all program states up to a bound, e.g., states where all integers are encoded using 8 bits)
on the equivalence of S and the translated program. VCs are logical expressions encoding the correctness properties of S. Specifically, given a program $P$ with $vars$ representing all the variables appearing in $P$, and $pre$, $post$, $inv$ representing the pre-conditions, the post-condition, and the invariant(s), respectively, the VCs for a program with loops consist of the following clauses:
@@
There exists standard techniques for generating VCs from a given source program [11]. In Tenspiler, the PS and invariants in the VCs are generated as placeholders as S is analyzed,
with their bodies to be synthesized during the synthesis phase.
@@
The goal of synthesis is then to find expressions from the search space $G$ for $PS$ and $inv$s such that, for all program states $\sigma$, they satisfy the VC. The synthesizer’s effectiveness in solving Eq. 2 depends on the search space’s description, as an overly expressive grammar can degrade synthesis performance. Conversely, a restrictive grammar may cause synthesis to fail in finding suitable expressions.
Various approaches exist for implementing the search to solve Eq. 2, from symbolic approaches like enumeration, deductive search, or constraint solving, to neural-based methods. For Tenspiler, we use an off-the-shelf symbolic search engine, Rosette [34], which uses constraint solving to address the synthesis problem. In a constraint solving-based approach, the specification $\phi$ (i.e., VC) and the search space $G$ are encoded as a single formula, and an SMT solver is then utilized to find a model that satisfies the formula. As a constraint-based solver, increasing the number of constraints makes the problem more challenging. Given that $\phi$ is fixed for a particular benchmark, the design of the $G$ becomes crucial. In Sec. 5, we discuss how the design of Tensir helps keep the grammar size reasonable and scales the synthesis process.
@@ 4.2 Verification
During the synthesis phase, as the PS and loop invariant(s) are validated against the VC only for a bounded set of program states, [^4] We are unaware of any SyGuS solvers that can validate against an unbounded set of program states efficiently, including state of the art solvers such as Z3 and CVC5. it is essential to check their validity for all program states. Tenspiler uses an SMT solver to do so by negating the VC in program verification i.e., checking if $\;\neg VC(S,\;inv_{1},\;inv_{2},\;...,\;PS,\;\sigma)$
is satisfiable for some $\sigma$. The placeholders in the VC are substituted with the synthesized bodies of $PS$ and $inv$s. If the solver cannot find any such $\sigma$, then the generated $PS$ and $inv$s are correct for all possible program states, thus proving $PS$ and $inv$s hold for all program states. If a $\sigma$ is found, then Tenspiler will iterate back to the synthesis phase in search of another candidate expression.
Besides using SMT solvers for Eq. 2, Tenspiler also leverages SMT solvers’ support of algebraic data types (ADT) to allow users to define common data structures such as lists and tuples. Internally, Tenspiler models tensors using the list data structure defined using ADTs. We use ADT’s accessor and constructor functions to retrieve and create new tensors. All the tensor accessing functions like slice, take are modeled as recursive functions over the list data structure.
Currently, while image processing kernels use integers and deep learning kernels operate over floats, we verify all the benchmarks using the theory of integers and reals, due to poor solver support for reasoning about floats.
@@ 4.3 Code Generation
After successful verifying the synthesized Tensir program, the final stage in Tenspiler’s workflow is to translate the Tensir program into the concrete syntax of a target DSL. Tensir makes this easy as it inherently represents tensor operations supported by all the target DSLs, and code can be generated using simple syntax-driven translation from Tensir operators to their DSL-specific counterparts.
1# Generating code for TPC-C language of Intel’s Gaudi processor
2gaudi_translations = {
3 "vec_elemwise_add": lambda args: f"v_u8_add_b({codegen(args[0],gaudi_translations)},{codegen(args[1],gaudi_translations)})",
4 ... }
5# Generating code for TensorFlow/PyTorch/MLX/NumPy DSLs
6tf_pt_mlx_np_translations = {
7 "vec_elemwise_add": lambda args: f"{codegen(args[0],tf_pt_mlx_np_translations)}+{codegen(args[1]),tf_pt_mlx_np_translations}",
8 ... }
9def codegen(expr: Expr, translations: Dict[str, Any]) -> str:
10 if isinstance(expr, CallExpr): return translations[expr.name()](expr.args)
@@
Fig. 5 shows part of Tenspiler’s code generator for vec_elemwise_add operator, which is a recursive function that takes as input a Tensir expression and translates every occurrence of vec_elemwise_add into the corresponding operation in the target DSL. For TPC-C [22], which is C-based programming language supported by Intel’s Gaudi processor, it generates v_u8_add_b intrinsic. [^5] To be precise, our code generator generates v_u8_add_b intrinsic for unsigned chars, while for float32 vectors it would generate v_f32_add_b. Meanwhile, it generates the + operator for TensorFlow, PyTorch, MLX, and NumPy.
This direct and syntactic translation simplifies the integration of new tensor-based target DSL into Tenspiler, as one would only need to add simple translation rules in the code generation process. For instance, we add support for MLX by adding only 65 lines of code to an existing 130-line template, as its API closely follows that of NumPy.
As a part of this work, we have implemented support for six different target DSLs in our code generator: NumPy, TensorFlow, PyTorch, MLX (an ML framework for Apple silicon), TPC-C (C-based programming language for Intel’s Gaudi processor), and Gemmini (an open-source neural network accelerator generator). We provide further details of these DSLs, including the sizes of all the benchmark programs translated in those DSLs by Tenspiler, in Appendix A.

### v3

# Tenspiler: A Verified-Lifting-Based Compiler for Tensor Operations

@@ Abstract
To enable developers in leveraging such DSLs while preserving their current programming paradigm, we present Tenspiler, a verified-lifting-based compiler that uses program synthesis to translate sequential programs written in general-purpose programming languages (e.g., C++ or Python code that does not leverage any specialized framework or accelerator) into tensor operations. Central to Tenspiler is our carefully crafted yet simple intermediate language, named Tensir, that expresses tensor operations. Tensir enables efficient lifting, verification, and code generation. Unlike classical pattern-matching-based compilers, Tenspiler uses program synthesis to translate input code into Tensir, which is then compiled to the target API / ISA. Currently, Tenspiler already supports six DSLs, spanning a broad spectrum of software and hardware environments. Furthermore, we show that new backends can be easily supported by Tenspiler by adding simple pattern-matching rules for Tensir. Using 10 real-world code benchmark suites, our experimental evaluation shows that by translating code to be executed on 6 different software frameworks and hardware devices, Tenspiler offers on average 105$\times$ kernel and 9.65$\times$ end-to-end execution time improvement over the fully-optimized sequential implementation of the same benchmarks.
@@ 1 Introduction
In this paper, we describe a tensor compiler that addresses these challenges. We introduce Tenspiler—a compiler designed to automate the transpilation of code to multiple tensor processing frameworks and hardware accelerators. Tenspiler uses verified lifting [12] (VL), a technique using inductive program synthesis to infer provably equivalent program summaries expressed using a user-defined intermediate representation (IR), and generate executable code from the synthesized summary to the target DSL. In contrast to conventional compilers that rely on pattern-matching to compile code, VL uses a search-based technique for the translation process. The two key steps of VL are:
@@
  • Verification: Once lifted, a theorem prover is used to validate if the synthesized summary is functionally equivalent to the input. If so, executable code is produced by calling the user-provided code generator from the summary; otherwise, another summary is generated by the search phase.
The key to making lifting efficient lies in the design of the IR (i.e., how the target DSL is modeled). In prior work [5, 6, 7, 34, 22], each function or instruction exposed by the target DSL is modeled explicitly. While doing so makes the search efficient, such explicit modeling makes the compiler hard to extend to other DSLs. With Tenspiler, we introduce, for the first time, a single unified IR, Tensir, that is designed for tensor operations and can easily generate code to multiple tensor processing software frameworks and hardware accelerators. Surprisingly, Tensir is a small language based on tensor algebra that includes commonly used vector and matrix operations. While other unifying IR exists (e.g., MLIR [24]), they are targeted for classical pattern-matching compilers. As we will discuss in Sec. 5 and Section 4.2, Tensir instead is designed for synthesis-based compilers and thus aims to enable both efficient search and verification.
In summary, this paper makes the following contributions:
@@
    We implement Tenspiler, a verified lifting-driven transpiler built using Tensir as the modeling language. We demonstrate the effectiveness of Tenspiler by using it to lift real-world code from 10 different suites to 6 different open-source and commercially-available tensor processing software frameworks and hardware accelerators. We illustrate the ease of constructing such transpilers by building one for MLX, a new tensor processing framework that was released only four months ago, using less than 200 lines of code in Sec. 6.
We have released Tenspiler’s code on https://github.com/tenspiler/tenspiler.
@@ 2 Overview
Figure 1: End-to-End example of using Tenspiler to transpile code.
Tenspiler takes in C/C++ or Python code as input [^1] Tenspiler currently supports a subset of the C/C++ and Python language (in particular it does not support code that uses pointers or objects, which we have not encountered such use in our benchmarks). It also expects any external libraries used in the input to be functionally modeled, which is how Tenspiler currently supports code that uses the STL::vector library. and transpiles it to a functionally equivalent program that leverages different software frameworks and hardware accelerators (details described in  Section 4.3) for tensor computation. As mentioned, Tenspiler uses verified lifting to first translate the source program into Tensir. Unlike traditional pattern-matching compilers, Tenspiler formulates code translation as a search for a program expressed in Tensir that is provably semantic-equivalent to the input. Doing so avoids the need to devise pattern-matching rules and prove their correctness. To make the search scalable, instead of directly searching within the DSL exposed by each target, we designed a high-level IR called Tensir that abstracts away the low-level implementation details of each DSL operator and captures only their semantics, unifying various DSLs into a common set of tensor operators. Tenspiler uses a program synthesizer (currently Rosette [36], a synthesizer for finite domain theories) to lift the input code to Tensir during the search phase. The synthesized output is then verified using a theorem prover (currently, an SMT solver, CVC5 [10]) for the unbounded domain. “Unbounded domain” means the verification is performed for all possible program states, not just a bounded set of states (e.g., all states where integers are represented using 8 bits) that Rosette considers during the synthesis phase. Once verified, Tenspiler then translates the Tensir program to the concrete syntax of the target DSL using simple pattern-matching rules.
@@
Tenspiler uses symbolic search to solve the synthesis problem. Symbolic search is typically implemented through enumerative or deductive search, and using constraint-solving approaches which often rely on domain-specific heuristics to scale. As a SyGuS problem, symbolic search is implemented as enumerating different expressions over a user-provided grammar, where the grammar encodes all possible combinations of operators in the target DSL up to a specified depth. However, as the depth increases, the number of choices grows exponentially, making the search intractable. As we will discuss in Sec. 5, Tensir is designed to make synthesis scalable. For S, the synthesis phase returns the following solution:
@@ 3 The Tensir Intermediate Representation
Comparison with MLIR. MLIR [24] is a compiler infrastructure that enables the representation and transformation of code at various levels of abstraction. The core idea behind MLIR is to provide a unified IR that can capture the semantics of the program at different levels of detail (dialects), from high-level abstractions down to low-level, target-specific instructions. Developers can use MLIR by progressively lowering the code through different dialects until it reaches a level suitable for the target hardware. While MLIR and its dialects offer a powerful infrastructure for progressively targeting multiple hardware backends, we found that the existing dialects do not fully support all the operators required for our use case. Independently, both the linalg and tensor dialects do not support all the operators Tensir supports. For example, the $select$ operator, which is crucial for image processing kernels that apply operations conditioned on pixel values, is not supported by any MLIR dialects. Additionally, unifying these dialects can be challenging for developers. Instead of unifying, recent work such as mlirSynth [14] has explored using program synthesis to translate between different MLIR dialects. In contrast, Tensir is designed to be flexible and easily extensible. Developers can add new operators to Tensir by simply describing their high-level semantics, and new backend support can be incorporated by defining simple pattern-matching rules. This approach allows developers to extend Tensir without going into the intricacies of MLIR. Moreover, Tensir can practically be compiled into different MLIR dialects, providing developers the flexibility to leverage the MLIR infrastructure if desired.
@@ 3.1 Language Definition
$a\in accessOp$ ::=:= $take(t,l)\;\mid\;tail(t,l)\;\mid\;slice(t,l,l_{1},l_{2})$
$l$ ::=:= $Integer\;Literal\;\mid\;size(t,l)\;\mid\;s_{comp}$
@@
Figure 2: Tensir grammar.
The recursive nature of Tensir’s grammar allows tensor operations to be composed, facilitating the expression of complex algorithms encountered in source code. Tensir extends its expressiveness beyond tensor operations by also including a control flow operator (controlflowOp). It integrates control flow through the ite operator, enabling conditional logic into tensor computation. This operator is crucial for translating real-world loopy programs that contain branches.
@@
  • slice(t, l, s, e) extracts a contiguous sub-tensor from t from indices s (inclusive) to e (exclusive) along dimension l.
  • size(t, l) returns the size of tensor t in dimension l.
@@ 4 Transpiling Code Using Tenspiler
As shown in Fig. 3, Tenspiler is designed to translate a program in high-level languages, source (S), into another program that leverages different tensor processing infrastructures. Tenspiler currently support a vector processing library (NumPy) for CPU execution, DL frameworks (PyTorch, TensorFlow, MLX) for GPU execution, and ISAs for specialized hardware accelerators (Gaudi, Gemmini). Tenspiler is a verified-lifting-based compiler, leveraging search to find a program within the target domain. Instead of relying on traditional pattern-matching rules, Tenspiler translates source programs with a 3-phase workflow: 1. synthesis, 2. verification, and 3. backend code generation. Tenspiler uses a single IR, Tensir, to facilitate all 3 phases of its workflow. Tensir is designed to include tensor processing operators common to all target backends. As shown in the figure, the synthesis phase takes in S and generates a program summary expressed using Tensir. Then, in the verification phase, Tenspiler verifies the generated summaries to ensure their semantic equivalence with S. Finally, in the code generation phase, the Tensir program is translated to the concrete syntax of the target DSL(s).
@@ 4.1 Synthesis
For Tenspiler, the specification is to find a functionally equivalent program to S. Various methods exist to express this specification, such as using input-output examples, bounded-model checking, and verification conditions (VC) [19].
In Tenspiler, we use VCs as the specification for the synthesis phase as it provides full guarantees (i.e., for all program states up to a bound, e.g., states where all integers are encoded using 8 bits) on the equivalence of S and the translated program. VCs are logical expressions encoding the correctness properties of S.
Specifically, given a program $P$ with $vars$ representing all the variables appearing in $P$, and $pre$, $post$, $inv$ representing the pre-conditions, the post-condition, and the invariant(s), respectively, the VCs for a program with loops consist of the following clauses:
@@
There exist standard techniques for generating VCs from a given source program [11]. In Tenspiler, the $PS$ and invariants in the VCs are generated as placeholders as S is analyzed, with their bodies to be synthesized during the synthesis phase.
@@
The goal of synthesis is then to find expressions from the search space $G$ for $PS$ and $inv$s such that, for all program states $\sigma$, they satisfy the VC.
For Tenspiler, we use an off-the-shelf symbolic search engine, Rosette [36], which uses constraint solving to address the synthesis problem. In a constraint-solving-based approach, the specification $\phi$ (i.e., VC) and the search space $G$ are encoded as a single formula, and an SMT solver is then utilized to find a model that satisfies the formula. As a constraint-based solver, increasing the number of constraints makes the problem more challenging. Given that $\phi$ is fixed for a particular benchmark, the design of the $G$ becomes crucial. In Sec. 5, we discuss how the design of Tensir helps keep the grammar size reasonable and scales the synthesis process.
@@ 4.2 Verification
During the synthesis phase, as the $PS$ and loop invariant(s) are validated against the VC only for a bounded set of program states, [^4] We are unaware of any SyGuS solvers that can validate against an unbounded set of program states efficiently, including state of the art solvers such as Z3 and CVC5. it is essential to check their validity for all program states. Tenspiler uses an SMT solver to do so by negating the VC in program verification i.e., checking if $\;\neg VC(S,\;inv_{1},\;inv_{2},\;...,\;PS,\;\sigma)$ is satisfiable for some $\sigma$. The placeholders in the VC are substituted with the synthesized bodies of $PS$ and $inv$s. If the solver cannot find any such $\sigma$, then the generated $PS$ and $inv$s are correct for all possible program states, thus proving $PS$ and $inv$s hold for all program states. If a $\sigma$ is found, then Tenspiler will iterate back to the synthesis phase in search of another candidate expression.
Besides using SMT solvers for Eq. 2, Tenspiler also leverages SMT solvers’ support of algebraic data types (ADT) to allow users to define common data structures such as lists and tuples. Internally, Tenspiler models tensors using the list data structure defined using ADTs. We use ADT’s accessor and constructor functions to retrieve and create new tensors. All the tensor accessing functions like slice, take are modeled as recursive functions over the list data structure. Currently, while image processing kernels use integers and deep learning kernels operate over floats, we verify all the benchmarks using the theory of integers and reals, due to poor solver support for reasoning about floats.
@@ 4.3 Code Generation
After successfully verifying the synthesized Tensir program, the final stage in Tenspiler’s workflow is to translate the Tensir program into the concrete syntax of a target DSL. Tensir makes this easy as it inherently represents tensor operations supported by all the target DSLs, and code can be generated using simple syntax-driven rules that map Tensir operators to their DSL-specific counterparts.
To translate the Tensir expression into an executable DSL program, our code generation step recursively processes each part of the Tensir expression. Fig. 5 illustrates a portion of the code generation function for PyTorch. The function maps Tensir variables to their names (line 3), literals to their values (line 5), and function calls to their PyTorch equivalents based on function signatures (lines 6-15).
Consider the running example in Figure 1(a), where the synthesized Tensir solution is t_t(t_t(b, a, +), t_s(t_t(b, a, *), 255, /), -). This expression represents a t_t function call with the - operator, which maps to torch.subtract as shown in line 13. Next, the codegen function is called recursively on the two arguments, t_t(b, a, +) and t_s(t_t(b, a, *), 255, /). This results in the final translated PyTorch expression as torch.subtract(torch.add(b, a), torch.divide(torch.add(b, a), 255)).
To extend support for a new backend, one simply needs to replace the DSL operator names in lines 11, 15, 17, and 19. For example, in MLX’s codegen, torch.add on line 11 would be replaced by mlx.core.add.
1def codegen(expr: Expr):
2 if isinstance(expr, Var):
3 return expr.name()
4 elif isinstance(expr, Lit):
5 return expr.val()
6 elif isinstance(expr, Call):
7 f_name, args = expr.name(), expr.arguments()
8 if f_name in {"t_t", "t_s"}:
9 op = args[-1]
10 if op == "+":
11 return f"torch.add({codegen(args[0])},{codegen(args[1])})"
12 #corresponding MLX return statement
13 #return f"mlx.core.add({codegen(args[0])},{codegen(args[1])})"
14 elif op == "-":
15 return f"torch.subtract({codegen(args[0])},{codegen(args[1])})"
16 elif op == "*":
17 return f"torch.multiply({codegen(args[0])},{codegen(args[1])})"
18 elif op == "/":
19 return f"torch.divide({codegen(args[0])},{codegen(args[1])})"
@@
This direct and syntactic translation simplifies the integration of new tensor-based target DSL into Tenspiler, as one would only need to add simple translation rules in the code generation process. For instance, we add support for MLX by changing only 65 lines of code to an existing 200-line template, as its API closely follows that of NumPy.
[…]